# Analysis of static and dynamic batching algorithms for graph neural networks

### v2

@@ 5 Implementation
There are several open-source software libraries to choose from to perform these experiments. We opt to use the Jraph library (Godwin* et al., 2020), which is built using the JAX library (Bradbury et al., 2018). The JAX library traces out computations into a computation graph that it uses to optimize/compile the code. This compiled auto-differentiation code is optimized to run fast, especially on GPU. To run our experiments, we extend Jraph to run the static-$2^{N}$, static-64, and static-constant algorithms, log necessary profiling information and run the models we have selected for our tests.
6 Compute hardware
As mentioned, the benefit of dynamic batching is that the memory allocated to a batch of graphs is always the same for each batch. This allows the user to compile the update functions only once and optimize the compiled binary to run on GPU. Note that with static-64 and static-$2^{N}$ batching, the update function can also be compiled, but needs to be recompiled for different multiples of 64 / powers of two (the static-constant algorithm does not have this problem). We time the training of models for different batching algorithms. We run experiments on a cluster that comprises two Intel Xeon IceLake Platinum 8360Y CPUs and 4 NVIDIA A100 GPUs connected via NVLINK. Our experiments use a single GPU node and a single CPU core. This is done so that we can understand the algorithms performance in a controlled setting. We compare these experiments to those done without a GPU and run only on a single Intel Xeon IceLake Platinum 8360Y CPU core.
@@ 7 Models
We evaluate three models to get insight into any model dependent effects on our profiling results. The models take in a graph, $G$, composed of nodes (or vertices) and edges $G(n,e)$. The nodes are represented by feature vectors $h_{n}^{t}$, where the subscript represents the specific node $n$, and the superscript $t$ represents the number of times the vector has been updated.
@@
The SchNet model (Schütt et al., 2018) is a graph convolutional NN. The node feature vectors are updated with a node update equation (or convolution) that convolves the feature vector of the node, $h_{i}^{t}$, and feature vectors in neighborhood, $N_{i}$, of the node, $i$. It uses a convolutional kernel, $W(\vec{r_{j}}-\vec{r_{i}})$, which depends on the euclidean distance between nodes:
@@
We also look at a message passing NN model with edge updates (MPEU)  (Jørgensen et al., 2018; Bechtel et al., 2023). Message passing NNs make use of a message function $M_{ij}^{t}$, which is defined for a pair of nodes, $i$ and $j$, and their corresponding edge $e_{ij}^{t}$.
@@
$\odot$ denotes element-wise multiplication and $\sigma$ represents the shifted soft plus function (Zhao et al., 2018).
Finally, we evaluate the algorithms on the equivariant PaiNN model (Schütt et al., 2021). This model differs from the MPEU and SchNet by learning not only scalar features, $s_{i}\in\mathbb{R}^{Fx1}$, for the representation of each atom but also equivariant features, $\vec{v_{i}}\in\mathbb{R}^{Fx3}$. The two representations are used to update each other. The scalar representation is updated with the following equation:
@@
$\Delta\vec{v_{i}^{u}}=a_{vv}(s_{i},\lVert V\vec{v}\rVert)U\vec{v},$ (9)
The three models were chosen, since the SchNet model is often used as a benchmark, the MPEU has shown better results on the AFLOW dataset in a benchmark study (Bechtel et al., 2023), and the PaiNN model is widely used (Kubečka et al., 2024). These models vary in size. The SchNet model is the smallest with 84,768 trainable parameters. The next largest is the PaiNN model with 177,568 trainable parameters, and the MPEU model contains 2,553,472.
8 Profiling results
To understand how the batching algorithms work in practice, we analyze the statistics of the number of nodes in a batch before padding occurs. The histograms for both batching algorithms are shown for the QM9 dataset in Fig. 2 for 10,000 batches with a batch size of 32. As stated above, the dynamic batching algorithm checks the node and edge padding target, as well as the number of graphs already in the batch, before adding a graph to the batch. For this dataset and batch size, the dynamic algorithm’s node padding target is 576. We can see that all dynamic batches have fewer than 576 nodes before padding. This results in a one-sided distribution, roughly a truncated Gaussian distribution, where the right-hand side is cutoff near the budget. In the dynamic case, the distribution of edges before padding shows an even clearer cutoff. This is due to the fact that, more often, the budget for edges is reached sooner than the budget for nodes. For static batching, however, no such check exists, and we observe a roughly Gaussian distribution of nodes. Similar effects are seen in the figure for the number of edges before padding.
Figure 2: Histograms of the number of nodes (top) and edges (bottom) in a batch of size 32 before padding for the static-$2^{N}$ and dynamic batching algorithms running on the QM9 datset.
We perform timing experiments on both algorithms for four batch sizes (16, 32, 64, 128) using both datasets and the three models for two million training steps. This number of training steps was chosen since it resulted in converged models  (Bechtel et al., 2023). We run each combination of batch size, algorithm, model, and dataset ten times to limit noise effects from the hardware in the profiling results. Each experiment reports the mean batching time, mean gradient update step time, and the sum of the two (combined time). Below, we discuss the results on GPU, while results on CPU are found in the Appendix.
Figure 3: Mean batching time (upper row), mean gradient-update time (middle row), and the mean combined time (bottom row) for varying batch sizes on the AFLOW (left) and QM9 (right) data using the SchNet model. For each datapoint, ten iterations of two million training steps are run.
According to Fig. 3, for the SchNet model, the static-$2^{N}$ and static-64 algorithms perform batching at near identical speeds. The dynamic batching, however, is clearly slower. All three algorithms show roughly linear scaling of the batching time with respect to the batch size, however, dynamic batching scales worse for larger batch sizes (i.e., has a steeper slope). This reflects the expectation that the time complexity of the batching algorithms is governed by the batch size. The dynamic algorithm is slower due to the added overhead from the bookkeeping of the padding targets. The batching results depend slightly on the dataset, but the rankings do not change and the scaling appears similar. The results for the PaiNN model are shown in Fig. 4, while the results for the MPEU model are shown in Fig. 10 in the Appendix.
Figure 4: Mean batching time (upper row), mean gradient-update time (middle row), and mean combined time (bottom row) for varying batch sizes on the AFLOW (left) and QM9 (right) data using the PaiNN model. For each datapoint, ten iterations of two million training steps are run.
For the SchNet and MPEU models, the timing results of the gradient-update step at small batch sizes are similar for all algorithms. For the PaiNN model, the static-$2^{N}$ algorithm has a slower gradient update for all batch sizes. For larger batch sizes, across all models and datasets, the dynamic batching algorithm performs the gradient update the fastest. The dynamic batching algorithm is likely the fastest, since the gradient update code is never recompiled. As seen in the histograms in Fig. 1, the number of nodes in a batch before padding is typically smaller for the dynamic algorithm. This is also true after padding. GNNs like the models we examine here, typically have a node update equation, which means that inference, and therefore the gradient update step, scales roughly linearly with the number of nodes. When the difference between the number of nodes is small between the dynamic and static-64 algorithm we do not expect much difference in the gradient update step time. The static-$2^{N}$ algorithm, however, typically adds a large number of padded nodes, which explains why the algorithm is typically slower than the dynamic algorithm in the gradient update step. The static-64 and dynamic algorithm show roughly linear scaling in the update time with the batch size. The static-$2^{N}$ algorithm shows poor relative performance especially for larger batch sizes. More work needs to be done to understand the static-$2^{N}$ scaling behavior.
The sum of the batching and gradient-update step mean timings, i.e., the mean combined time, roughly gives us the mean time per training step. For the SchNet model, the static-$2^{N}$ model performs the combined steps the fastest for the AFLOW dataset and for the batch size 16 in the QM9 dataset. For the batch size 64 and 128, the static-64 algorithm runs fastest. For the MPEU model the static-$2^{N}$ algorithm is fastest for batch sizes 16, 32 and 64 on AFLOW but the static-$64$ algorithm is faster for the batch size 128 on AFLOW and for all batch sizes on QM9 data. The poor update step scaling of the static-$2^{N}$ algorithm becomes the limiting factor for larger batch sizes on the QM9 dataset. For the PaiNN model, the dynamic algorithm has the shortest combined time for the AFLOW dataset and has roughly the same time requirements as the static-64 algorithm for the QM9 dataset despite having the slowest batching performance. Since the PaiNN model architecture requires a much larger update time the relative savings of the dynamic algorithm in the gradient step outweigh the poorer batching performance.
We visualize the speedup from switching algorithms for the PaiNN model on the QM9 and AFLOW datasets in Fig. 5. The speedup is defined as the ratio between the slowest algorithm combined time per step, $T_{s}$, and the faster algorithm combined time per step, $T_{x}$, as $\frac{T_{s}}{T_{x}}$ (Hennessy & Patterson, 2011). For the PaiNN model and batch size 128, the dynamic batching and static-64 algorithm have a speedup of over 2.7 when compared to the the static-$2^{N}$ algorithm. Similarly, for the PaiNN model for AFLOW and batch sizes of 32, 64 and 128, the dynamic algorithm has a speedup of roughly 1.5x when compared to the static-$2^{N}$ algorithm. For the MPEU model on the QM9 dataset, we observe 1.5x speedup when using the static-64 algorithm compared to the static-$2N$ algorithm. These observations over different datasets and models point to the large effect that changing the batching algorithm can have on the training time.
Figure 5: Speedup when switching from the slowest algorithm in terms of combined training time per step for the PaiNN model for the batch size of 64 for the AFLOW (top) and QM9 (bottom) datasets. In this case the slowest algorithm is the static-$2^{N}$ algorithm and therefore the speedup is always one.
For fewer training steps, the number of recompilations required in the gradient-update step is the dominant factor defining the ranking of the algorithms. This number of recompilations is shown in Fig. 6 for the QM9 data for two million training steps. The number of recompilations is model and hardware independent. We see that the static-64 algorithm performs several orders of magnitude more recompilations than the static-$2^{N}$ and dynamic algorithms. This is to be expected, since the dynamic algorithm compiles only once, and if it encounters a graph, which is bigger than the budget, the program terminates. Moreover, for static-64 batching, JAX recompiles the gradient-update step function every time a new nearest multiple of 64 is encountered when padding, and for static-$2^{N}$ batching every time a new nearest power of two is encountered, which is less likely for large powers of two. Note that most of the recompilations for the QM9 dataset happen in the first hundred thousand training steps. For example, for the batch size 32, the static-$2^{N}$ algorithm recompiles four times while the static-64 algorithm recompiles 89 times. From training step one hundred thousand to two million, the static-$2^{N}$ algorithm does not recompile again and the static-64 algorithm recompiles only an additional twenty times. The effect of recompilations is seen in Fig. 9 and Fig. 8 in the Appendix which show the mean and median combined times after running only one hundred thousand training steps. In the figures, the static-64 algorithm has the lowest median update times while the static-$2^{N}$ has the lowest mean update times. The mean is affected by outliers while the median is not. This tells us that when removing the effect of outliers, i.e., slow gradient-update steps which are caused by recompilations, the static-64 algorithm is fastest to update the gradients. Further, this tells us that the algorithm training time rankings are dependent on the dataset (whose variability in graph sizes cause recompilations) and the of number training steps run.
Figure 6: Number of recompilations on the QM9 dataset after two million training steps in the gradient-update step as a function of batch size for different batching algorithms. Note, that these numbers are model and hardware independent.
The batching algorithm may also affect the learning of the model. Dynamic batching does not always include the same number of graphs in a batch, since one or more fake padded graphs are added when the node or edge budget is reached. These padded graphs do not contribute to the gradient in the update step. The histogram in Fig. 7 depicts the distribution of the number of graphs before padding for the AFLOW and QM9 dataset after running 10,000 iterations with the dynamic algorithm. While most batches contain 31 graphs for a batch size of 32, some batches for the AFLOW dataset contain as few as ten graphs from the training data. That the QM9 graph node and edge distribution have shorter tails than the AFLOW distribution, as seen in Fig. 1, may explain why the distribution of the number of graphs before padding has shorter tails for QM9 than for AFLOW.
Figure 7: Histogram of number of graphs before padding in a batch of size 32 for the AFLOW and QM9 dataset.
The effect of the algorithm on learning can be discerned by training each model to convergence. The test loss (RMSE) as a function of the number of training steps on the QM9 data is shown in the Appendix Fig. 14. Visually, we see general agreement across models but for some batch sizes there are significant discrepancies. To quantify the difference in test metrics between two batching methods, we perform a pairwise student t-test across the final RMSE test loss. Fig. 16 in the Appendix shows a heatmap of pairwise t-test values and Fig. 17 shows the associated p-values across batching methods for the SchNet model for the QM9 dataset and batch size 64. The complete set of plots for the other combinations of datasets, models and batch sizes can be found in the Data Availability section. One can either use the t-test statistic or the p-value to determine whether to reject the null hypothesis (Hartmann et al., 2023), i.e., that the test metrics after two million training steps from different batching methods belong to the same underlying distribution. We adopt a p-value criteria from the literature (Ross, 2017), that argues that there is little evidence that the distribution of the test metrics differ significantly if the p-values are greater than 0.10. For p-values between 0.05 to 0.10, there is moderate evidence that the distributions are different. For lower p-values the evidence is strong. Ignoring the p-values that are greater than 0.10, we can look at several interesting results. For the QM9 data, SchNet model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.033. For the QM9 data, SchNet model, and batch size 32, the static-$2^{N}$ and static-64 algorithms have p-values with the dynamic algorithm of 0.067 and 0.053, respectively. For QM9 data, MPEU model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.071. These results show that for some datasets, models, and batch sizes, the dynamic algorithm gives significantly different test metrics than the static algorithm. For the majority of combinations, however, the results are similar, which may be due to use of the Adam optimizer whose adaptive estimation of first and second moments may reduce the effect of occasionally having smaller batch sizes.
The static-constant algorithm was also evaluated for a subset of data, models and batch sizes. The results are shown in the Appendix Section F. The static-constant model performs poorly compared to the other static algorithms, especially for the AFLOW dataset. This is because the AFLOW dataset contains a long tailed distribution in terms of the node/edges and as a result the static-constant model uses a large padding target that slows down the timing results. As a result, we did not employ this method for the full set of data combinations in our timing experiments.
9 Discussion
Our results show that the timing performance of the batching algorithms is dataset, model, and batch size dependent. This is true for experiments on CPU and GPU. For some cases, choosing one algorithm versus another can result in more than a 2.7x speedup in training time. The findings show that users should, for every new dataset and model, perform an initial performance test of batching algorithms. For larger hyperparameter searches or neural architecture searches, this speedup can save a very large amount of computational resources.
Diving deeper into the timing performance, the static algorithms perform the batching step faster. This is due to the added overhead from the bookkeeping of the padding targets in the dynamic algorithm. All of the algorithms show linear scaling in the batching time with respect to the batch size. The static-64 and dynamic algorithms show linear scaling in the gradient-update step, while the static-$2^{N}$ algorithm shows exponential scaling. This is likely due to the exponentially larger memory sizes required when using powers of two for padding. The gradient-update step is performed fastest by the dynamic algorithm. This is partially because the algorithm is never recompiled but also because performing inference on padded graphs, which do not influence the loss function, is typically faster. Moreover, the the static-$2^{N}$ algorithm typically has a larger number of nodes in the batch, which slows down the gradient update step.
Software and Data
The software to perform the batching, profiling experiments, parsing of experiments, and plotting is found in https://github.com/speckhard/gnn_batching/. The student t-test heatmaps can be found at https://drive.google.com/drive/folders/1W4Fc0a7ne0z5vUDYgNc1_YuUNcjxwX3W?usp=sharing_.
Impact Statement
This paper presents work whose goal is to analyze and profile batching algorithms methods that reduce the environmental impact of training graph neural network models.
Appendix A Batch and pad algorithms

### v3

@@ 5 Implementation
There are several open-source software libraries to choose from to perform these experiments. We opt to use the Jraph library [jraph2020github], which is built using the JAX library [jax2018github]. The JAX library traces out computations into a computation graph that it uses to optimize and compile the code. This compiled auto-differentiation code is optimized to run fast, especially on GPU. To run our experiments, we extend Jraph to run the $2^{N}, static$-64, and static-constant algorithms, log necessary profiling information and run the models we have selected for our tests.
@@ 6 Models
We evaluate three models to get insight into any model dependent effects on our profiling results. Specifically, the SchNet model [schutt2018schnet], a message passing model with edge updates (MPEU) model [jorgensen2018neural, bechtel2023band] and the PaiNN model [schutt2021equivariant] which uses equivariant features. The three models were chosen, since the SchNet model is often used as a benchmark, the MPEU has shown better results on the AFLOW dataset in a benchmark study [bechtel2023band], and the PaiNN model is widely used [kubevcka2024accurate]. These models vary in size. The SchNet model is the smallest with 84,768 trainable parameters. The next largest is the PaiNN model with 177,568 trainable parameters, and the MPEU model contains 2,553,472.
7 Test setup and compute hardware
As mentioned, the benefit of dynamic batching and static-constant batching is that the memory allocated to a batch of graphs is always the same for each batch. This allows the user to compile the update functions only once and optimize the compiled binary to run on GPU. Note that with static-64 and static-$2^{N}$ batching, the update function can also be compiled, but needs to be recompiled for different multiples of 64 / powers of two (the static-constant algorithm does not have this problem). We time the training of models for different batching algorithms.
Batching and weight update timing experimental setup. We perform timing experiments on both algorithms for four batch sizes (16, 32, 64, 128) using both datasets and the three models for two million training steps. This number of training steps was chosen since it resulted in converged models  [bechtel2023band]. Experiments with fewer steps are run and the results can be found in the Appendix. We run each combination of batch size, algorithm, model, and dataset ten times to limit noise effects from the hardware in the profiling results. Each experiment reports the mean batching time, mean gradient update step time (update time), and the sum of the two (combined time).
Hardware used. We run experiments on a cluster that comprises two Intel Xeon IceLake Platinum 8360Y CPUs and 4 NVIDIA A100 GPUs connected via NVLINK. We ran experiments with using a single CPU core and single GPU. This is done so that we can understand the algorithms performance in a controlled setting. Results from running only on CPU and multi-GPU setups are described in the appendix.
8 Profiling results
Batch statistics To understand how the batching algorithms work in practice, we analyze the statistics of the number of nodes in a batch before padding occurs. The histograms for both batching algorithms are shown for the QM9 dataset in Fig. 1 for 10,000 batches with a batch size of 32. As stated above, the dynamic batching algorithm checks the node and edge padding target, as well as the number of graphs already in the batch, before adding a graph to the batch. For this dataset and batch size, the dynamic algorithm’s node padding target is 576. We can see that all dynamic batches have fewer than 576 nodes before padding. This results in a one-sided distribution, roughly a truncated Gaussian distribution, where the right-hand side is cutoff near the budget. For static batching, however, no such check exists, and we observe a roughly Gaussian distribution of nodes. Similar effects are seen in the figure for the number of edges before padding.
Figure 2: The running average of the combined time (sum of the batching step and gradient-update step and) required per training step as a function of the total number of training steps run. Here only a single iteration is run for the batch size 32, MPEU model and QM9 dataset for both the dynamic, static-64 and static-$2^{N}$ algorithms.
Recompilation. For fewer training steps, the number of recompilations required in the gradient-update step can be the dominant factor defining the ranking of the algorithms. The number of recompilations is shown in Fig. 3 for the QM9 data for two million training steps. The number of recompilations is model and hardware independent. We see that the static-64 algorithm performs several orders of magnitude more recompilations than the static-$2^{N}$ and dynamic algorithms. This is to be expected, since the dynamic algorithm compiles only once, and if it encounters a graph, which is bigger than the budget, the program terminates and needs to be restarted with a larger pardding target. Moreover, for static-64 batching, JAX recompiles the gradient-update step function every time a new nearest multiple of 64 is encountered when padding, and for static-$2^{N}$ batching every time a new nearest power of two is encountered, which is less likely for larger powers of two. For example, for the batch size of 32, the static-$2^{N}$ algorithm recompiles four times while the static-64 algorithm recompiles 89 times.
Most of the recompilations for the QM9 dataset happen in the first hundred thousand training steps. From training step one hundred thousand to two million, the static-$2^{N}$ algorithm does not recompile again and the static-64 algorithm recompiles only an additional twenty times. The effect of recompilations is seen in Fig. 6 and Fig. 7 in the Appendix which show the mean and median combined times after running only one hundred thousand training steps respectively. In the figures, the static-64 algorithm has the lowest median update times while the static-$2^{N}$ has the lowest mean update times. The mean is affected by outliers while the median is not. This tells us that when removing the effect of outliers, i.e., slow gradient-update steps which are caused by recompilations, the static-64 algorithm is fastest to update the gradients. This tells us that the algorithm training time rankings are dependent on the dataset (where more variability in graph sizes causes more recompilations) and the number of training steps run.
Figure 3: Left: number of recompilations on the QM9 dataset after two million training steps in the gradient-update step as a function of batch size for different batching algorithms. Note, that these numbers are model and hardware independent. Right: histogram of the number of graphs before padding in a batch size of 32 for the AFLOW and QM9 dataset.
Static-constant batching results. The static-constant algorithm was also evaluated for a subset of data, models and batch sizes. The results are shown in the Appendix Section H. The static-constant model performs poorly compared to the other static algorithms. As shown in Figure 1, the AFLOW dataset contains a long-tailed distribution in terms of the node/edges (the estimated excess kurtosis of the two histograms are greater than one [decarlo1997meaning]). Moreover, the mean number of nodes/edges is 10-20 times smaller than the maximum number. As a result, for AFLOW, the static-constant model uses a large padding target that slows down the batching and update step. For QM9, the histograms of the dataset show that the distribution of nodes and edges resemble more closely a normal distribution. The number of edges varies from roughly one hundred to six hundred edges, and as a result, for the QM9 data, we see a smaller yet still significant slowdown using the static-constant batching algorithm compared to other methods.
Static-64, static-$2^{N}$ and dynamic algorithm results Fig. 4 displays the breakdown of batching time per training step. We see the batching times appear roughly model and dataset independent and scale linearly with the batch size. The static algorithms loop over the number of graphs in the batch, and the dynamic algorithm does the same but may exit the loop early based on the padding target checks. Therefore it is not surprising that we the batching steps scale $\mathcal{O}(n)$ where n is the batch size. The dynamic batching is slower than the static-$2^{N}$ and static-64 algorithms and is due to the added overhead from the bookkeeping of the padding targets.
Figure 4: Timing measurements while varying the batch size for SchNet (left two columns), MPEU (middle columns) and PaiNN (right two columns) Results for the AFLOW and QM9 dataset are shown for each model. For each datapoint, ten iterations of two million training steps are run.
For the SchNet and MPEU models, the timing results of the gradient-update step at small batch sizes are similar for all algorithms. For the PaiNN model, the static-$2^{N}$ algorithm has a slower gradient update for all batch sizes. For larger batch sizes, across all models and datasets, the dynamic batching algorithm performs the gradient update the fastest. The dynamic batching algorithm is the fastest, since the gradient update code is never recompiled. As seen in the histograms in Fig. 1, the number of nodes in a batch before padding is typically smaller for the dynamic algorithm. This is also true after padding. GNNs like the models we examine here, typically have a node update equation, which means that inference, and therefore the gradient update step, scales roughly linearly with the number of nodes. When the difference between the number of nodes is small between the dynamic and static-64 algorithm we do not expect much difference in the gradient update step time. The static-$2^{N}$ algorithm, however, typically adds a large number of padded nodes, which explains why the algorithm is typically slower than the dynamic algorithm in the gradient update step. The static-64 and dynamic algorithm show roughly linear scaling in the update time with the batch size. The static-$2^{N}$ algorithm shows poor relative performance especially for larger batch sizes. More work needs to be done to understand the static-$2^{N}$ scaling behavior.
The sum of the batching and gradient-update step mean timings, i.e., the mean combined time, gives us the mean time per training step. The speedup is defined as the ratio between the slowest algorithm combined time per step, $T_{s}$, and the faster algorithm combined time per step, $T_{x}$, as $\frac{T_{s}}{T_{x}}$ [hennessy2011computer]. For the PaiNN model, QM9 dataset, and batch size 128, the dynamic batching and static-64 algorithm have a speedup of over 2.7 when compared to the static-$2^{N}$ algorithm. Similarly, for the PaiNN model for AFLOW and batch sizes of 32, 64 and 128, the dynamic algorithm has a speedup of roughly 1.5x when compared to the static-$2^{N}$ algorithm. For the MPEU model on the QM9 dataset, we observe 1.5x speedup when using the static-64 algorithm compared to the static-$2N$ algorithm. These observations over different datasets and models point to the large effect that changing the batching algorithm can have on the training time. The speedup from switching algorithms for the PaiNN model on the QM9 and AFLOW datasets can be seen in the Fig. 5. If we include the static-constant algorithm, which is always the slowest algorithm for the parameter space we explored here, the largest speedup we see is 12.5 for the PaiNN model on AFLOW with a batch size of 16, when switching to the dynamic algorithm from the static-constant algorithm.
Effect on learning. The batching algorithm may also affect the learning of the model. Dynamic batching does not always include the same number of graphs in a batch, since one or more fake padded graphs are added when the node or edge budget is reached. These padded graphs do not contribute to the gradient in the update step. The histogram in Fig. 3 depicts the distribution of the number of graphs before padding for the AFLOW and QM9 dataset after running 10,000 iterations with the dynamic algorithm. While most batches contain 31 graphs for a batch size of 32, some batches for the AFLOW dataset contain as few as ten graphs from the training data. The fact that the QM9 graph node and edge distribution have shorter tails than the AFLOW distribution, as seen in Fig. 1, may explain why the distribution of the number of graphs before padding has shorter tails for QM9 than for AFLOW. The effect of the algorithm on learning can be discerned by training each model to convergence. The test loss (RMSE) as a function of the number of training steps on the QM9 data is shown in the Appendix Fig. 12. Visually, we see general agreement across models but for some batch sizes there are significant discrepancies.
To quantify the difference in test metrics between two batching methods, we perform a pairwise Student’s t-test across the final RMSE test loss. Fig. 13 in the Appendix shows a heatmap of pairwise t-test values and Fig. 14 shows the associated p-values across batching methods for the SchNet model for the QM9 dataset and batch size 64. The complete set of plots for the other combinations of datasets, models and batch sizes can be found in the Data Availability section.
Figure 5: Speedup, on GPU, when switching from the slowest algorithm in terms of combined training time per step for the PaiNN model (not including the static-constant model) for the AFLOW (top) and QM9 (bottom) datasets. For both datasets the slowest algorithm is the static-$2^{N}$ algorithm. If the static-constant algorithm is included the speedup increases to a maximum of to 12.5
One can either use the t-test statistic (critical value approach) or the associated p-value to determine whether to reject the null hypothesis [hartmann2023stats] that the test metrics after two million training steps from different batching methods belong to the same underlying distribution. We adopt a p-value criteria from the literature [ross2017introductory], that argues that there is little evidence that the distribution of the test metrics differ significantly if the p-values are greater than 0.10. For p-values between 0.05 to 0.10, there is moderate evidence that the distributions are different. For lower p-values the evidence is strong. Ignoring the p-values that are greater than 0.10, we can look at several interesting results. For the QM9 data, SchNet model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.033. For the QM9 data, SchNet model, and batch size 32, the static-$2^{N}$ and static-64 algorithms have p-values with the dynamic algorithm of 0.067 and 0.053, respectively. For QM9 data, MPEU model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.071. These results show that for some datasets, models, and batch sizes, the dynamic algorithm gives significantly different test metrics than the static algorithm. For the majority of combinations, however, the results are similar, which may be due to use of the Adam optimizer whose adaptive estimation of first and second moments may reduce the effect of occasionally having smaller batch sizes.
9 Discussion
We find the static-constant algorithm to perform the worst in terms of timing performance per training step. We do not recommend using the algorithm unless the user has a dataset where the maximum graph size is very similar to the mean graph size, and/or the priority is to have robustness in the training pipeline. We see in some cases a speedup of up to 12.5x when switching from the static constant algorithm to the other algorithms. If we discard the static-constant algorithm, choosing one algorithm versus another can result in more than a 2.7x speedup in training time without observable differences in learning metrics. For larger hyperparameter searches or neural architecture searches (NAS), this speedup can save a very large amount of computational resources. Across, the parameter space explored in this work, the static-64 and dynamic algorithm appear the fastest. In general, we would recommend the dynamic algorithm since it performs faster than the static-64 algorithm for fewer number of training steps, which can be important for NAS.
Diving deeper into the timing performance, the static algorithms perform the batching step faster. This is due to the added overhead from the bookkeeping of the padding targets in the dynamic algorithm. All of the algorithms show linear scaling in the batching time with respect to the batch size. The static-64 and dynamic algorithms show linear scaling in the gradient-update step, while the static-$2^{N}$ algorithm shows exponential scaling. This is due to the exponentially larger memory sizes required when using powers of two for padding. The gradient-update step is performed fastest by the dynamic algorithm. This is partially because the algorithm is never recompiled. Moreover, the static-$2^{N}$ algorithm typically has a larger number of nodes/edges in the batch, which slows down the gradient update step.
Software and data availability
The software to perform the batching, profiling experiments, parsing of experiments, and plotting is found in https://github.com/speckhard/gnn_batching/. The complete set of Student’s t-test heatmaps can be found at https://drive.google.com/drive/folders/1W4Fc0a7ne0z5vUDYgNc1_YuUNcjxwX3W?usp=sharing_.
Appendix A Model descriptions
@@
In this section we describe the three models used to evaluate the batching algorithms in greater detail. The models take in a graph, $G$, composed of nodes (or vertices) and edges $G(n,e)$ [liu2022introduction, koller2009probabilistic, speckhard2025grid]. The nodes are represented by feature vectors $h_{n}^{t}$, where the subscript represents the specific node $n$, and the superscript $t$ represents the number of times the vector has been updated.
@@
The SchNet model [schutt2018schnet] is a graph convolutional NN. The node feature vectors are updated with a node update equation (or convolution) that convolves the feature vector of the node, $h_{i}^{t}$, and feature vectors in neighborhood, $N_{i}$, of the node, $i$. It uses a convolutional kernel, $W(\vec{r_{j}}-\vec{r_{i}})$, which depends on the euclidean distance between nodes:
@@
We also look at a message passing NN model with edge updates (MPEU)  [jorgensen2018neural, bechtel2023band]. Message passing NNs make use of a message function $M_{ij}^{t}$, which is defined for a pair of nodes, $i$ and $j$, and their corresponding edge $e_{ij}^{t}$.
@@
$\odot$ denotes element-wise multiplication and $\sigma$ represents the shifted soft plus function [zhao2018novel].
Finally, we evaluate the algorithms on the equivariant PaiNN model [schutt2021equivariant]. This model differs from the MPEU and SchNet by learning not only scalar features, $s_{i}\in\mathbb{R}^{Fx1}$, for the representation of each atom but also equivariant features, $\vec{v_{i}}\in\mathbb{R}^{Fx3}$. The two representations are used to update each other. The scalar representation is updated with the following equation:
@@
$\Delta\vec{v}_{i}^{u}=a_{vv}(s_{i},\lVert V\vec{v}\rVert)U\vec{v}$ (7)
Appendix B Batch and pad algorithms